# Logical computation demonstrated with a neutral atom quantum processor

### v2

@@ Abstract
Transitioning from quantum computation on physical qubits to quantum computation on encoded, logical qubits can improve the error rate of operations, and will be essential for realizing valuable quantum computational advantages. Using a neutral atom quantum processor with $256$ qubits, each an individual Ytterbium atom, we demonstrate the entanglement of $24$ logical qubits using the distance-two $\llbracket 4,2,2\rrbracket$ code, simultaneously detecting errors and correcting for lost qubits. We also implement the Bernstein-Vazirani algorithm with up to $28$ logical qubits encoded in the $\llbracket 4,1,2\rrbracket$ code, showing better-than-physical error rates. We demonstrate fault-tolerant quantum computation in our approach, guided by the proposal of Gottesman (2016), by performing repeated loss correction for both structured and random circuits encoded in the $\llbracket 4,2,2\rrbracket$ code. Finally, since distance-two codes can correct qubit loss, but not other errors, we show repeated loss and error correction using the distance-three $\llbracket 9,1,3\rrbracket$ Bacon-Shor code. These results begin to clear a path for achieving scientific quantum advantage with a programmable neutral atom quantum processor.
I Introduction
Quantum computers can enable faster, more accurate solutions than classical computers. However, the biggest challenge to scaling up quantum computers is errors, or “noise.” Physical qubits and their operation are inherently noisy. State-of-the-art two-qubit gate error rates depend on the qubit implementation, and in some flagship quantum processors range, e.g., from $p=0.1\%$ for an ion trap processor MBA+ (23) to an average of $p=0.4\%$ for a superconducting qubit processor GC (24). If a circuit cannot tolerate errors, it can only run for roughly $1/p$ steps, or fewer, which is not nearly enough to achieve large-scale, advantageous computation. Even the best physical qubits are insufficiently reliable, without error correction, to run even moderate-size quantum algorithms.
Indeed, obtaining a quantum advantage in scientific applications will likely require at least $100$ qubits with error rate $p$ at most $10^{-6}$, while a quantum advantage for industrial-scale problems may require upwards of $1000$ qubits with $p\sim 10^{-12}$ BMT+ (22). Such error rates will not be achievable at the physical level. Fortunately, encoding qubits in quantum error-correcting codes, thereby creating “logical” qubits, should allow dramatic reductions in effective error rates. The operations must also be encoded, according to fault-tolerant constructions AGP (06). Thus, by using quantum error correction and fault tolerance, noise in quantum hardware can be addressed through software, albeit with a cost in physical qubit overhead.
In this work, we present a step toward achieving scientific quantum advantage by demonstrating reliable quantum computation using logical qubits that outperforms corresponding physical-qubit based computation. We present our results on a neutral atom quantum processor. Neutral atom quantum architectures recently have been in rapid development. Previously, neutral atom architectures were predominantly analog in nature. Recent progress has enabled neutral atom approaches to transition to digital, gate-based computation EBK+ (23); SST+ (23); MLP+ (23); RCC+ (24); FTS+ (24); MSS+ (24). The promise of neutral atoms as qubits is their configurability, both in count and connectivity, while also maintaining low error rates. The neutral atom quantum processor we use in this work has up to 256 qubits (Ytterbium atoms), with all-to-all connectivity enabled by atom movement, along with high two-qubit physical gate fidelity and atom loss detection. This allows for flexible computation, as well as the ability to both use and optimize over a wide range of quantum error-correcting codes, including those requiring non-local connectivity, for the underlying neutral atom architecture. We refer to the system for detecting and correcting errors, also during computation, as the qubit virtualization system, as encoded, logical qubits are virtual qubits made up of some number of physical qubits.
Through co-optimization of fault-tolerant quantum error correction protocols within the qubit virtualization system and the neutral atom quantum processor design, we demonstrate the creation of logical qubits, their logical entanglement, and logical computation, while correcting for lost atoms.
Figure 1: Summary of results for the $24$-qubit cat state. Shown is the error for measurements in the $X$ basis, the error for measurements in the $Z$ basis, and their sum, for three cases. A total less than $50\%$ implies entanglement. The baseline (left) is an unencoded $24$-qubit cat state; the error bars are large because we have conservatively chosen only the best run instead of averaging all baseline runs, which would further increase the separation between logical and physical performance. See Fig. 7 for more details. The encoded bars are based on different ways of interpreting the same data for $\llbracket 4,2,2\rrbracket$-encoded cat states. In one case (middle), we reject trials with a detected error, but attempt to correct for lost qubits. In the other case (right), we reject trials with any lost qubit or detected error.
@@ I Introduction
Figure 2: Summary of results for the Bernstein-Vazirani algorithm with an $n$-bit secret string. The string is always $1^{n}$, as the corresponding oracle is the most difficult to implement. The success probability is the probability of measuring $1^{n}$, conditioned on acceptance. For the unencoded algorithm (black), we accept when the $n$ measured atoms are present. For the encoded algorithm (red), we accept when all of the $n$ measured code blocks are decodable; in particular, it can detect faults and correct for lost atoms. The lighter points are based only on those trials with no initial atom loss.
Key results. We demonstrate a variety of reliable quantum computations that outperform physical qubit computation using logical qubits co-designed with a programmable neutral atom quantum processor. First, we show entanglement of $24$ logical qubits in a “cat” state (Fig. 1). This is the largest encoded cat state on record to date, illustrating the scale available in neutral atom processors. We then demonstrate quantum algorithms encoded into logical qubit computations. We implement the Bernstein-Vazirani algorithm BV (97); CEMM (98), a quantum algorithm well known for showing the higher efficiency of quantum computers versus classical. Our implementations involve as many as $28$ logical qubits, encoded in $112$ physical qubits (Fig. 2). Next, we demonstrate fault tolerance guided by a proposal of Gottesman Got (16). These experiments incorporate repeated error detection and, notably, repeated loss correction, a critical feature for scaling neutral atom processors. Each of the above experiments are encoded to distance two; they are capable of detecting phase and bit-flip errors and correcting atom loss. Finally, we demonstrate correcting both loss and errors, with repeated error correction in the distance-three Bacon-Shor code Bac (06).
@@
Related work. Creating and verifying large cat states, also known as GHZ states GHZ (89), is a common benchmark for measuring hardware progress Kre (24). The quality of an experiment is typically expressed by the fidelity with the ideal cat state, and where a fidelity greater than $1/2$ is the success criterion. For example, Moses et al. MBA+ (23) prepared the $20$-qubit cat state with an $86\%$ fidelity, and the $32$-qubit cat state with an $82\%$ fidelity, on the Quantinuum H2 trapped-ion processor. Using superconducting qubits, Bao et al. BXS+ (24) prepared a $60$-qubit cat state with a $59\%$ fidelity. For encoded qubits, the published record to date is an encoded $12$-qubit cat state with at least a $99.8\%$ fidelity, also on H2 RAC+ (24). On a neutral atom system, Bluvstein et al. BEG+ (24) have prepared a four-qubit encoded cat state with a fidelity of $72(2)\%$ using error correction, ranging to $99.85^{+0.1}_{-1.0}\%$ using full error detection and therefore with much lower acceptance probability. Table 1 summarizes results with encoded cat states.
@@
Reference qubits Code Fidelity
HDHL (24) 4 $\llbracket 25,4,3\rrbracket$ $99.5^{+0.2}_{-0.4}\%$ to $99.7^{+0.2}_{-0.3}\%$
@@
72(2)error correction
$99.85^{+0.1}_{-1.0}\%$ error detection
RAC+ (24) 12 $\llbracket 16,4,4\rrbracket$ $99.82^{+0.12}_{-0.4}\%$ to $99.90^{+0.1}_{-0.3}\%$
@@
$89.8^{+4.7}_{-6.6}\%$ to $91.5^{+4.3}_{-6.2}\%$
error and loss detection
@@
The Bernstein-Vazirani algorithm for finding a secret string is another benchmark for quantum processors LGM+ (24). It has been implemented with unencoded qubits on superconducting processors RHK+ (20); DH (21); PL (23), ion trap processors FHM+ (16); DLF+ (16); WBD+ (19), and both processor types LMRM (17); MLM+ (19); BWM (21); LJV+ (23), for various lengths of the secret string, up to length $26$ for superconducting qubits PL (23) and up to length $20$ for ion traps LJV+ (23). It is difficult to compare these results fairly. Some test identification of the secret $s=1^{n}$, since it is hardest to implement; the corresponding oracle requires the most two-qubit gates. Others test against random strings $s$ or average over different Hamming weights, and this is much easier, particularly on ion traps, comparable to finding $s=1^{n/2}$. The success probabilities also vary considerably, e.g., from an exponentially small value for random $s$ with length $26$ PL (23), to $70\%$ for random $s$ with length $20$ LJV+ (23).
A critical capability for enabling more complex, reliable computation in a quantum processor is repeated detection and correction of errors. The latest generations of quantum hardware have enabled a variety of such demonstrations within the past few years. Repeated error detection with distance-two codes has been demonstrated on ion traps, including three rounds on eight logical qubits SBA (24), and 15 rounds on four logical qubits YDKR (24). Demonstrations based on Gottesman’s protocol Got (16) for fault tolerance with the $\llbracket 4,2,2\rrbracket$ code, but without including repeated error detection, include WWJ+ (18); HF (19); Vui (18); LGL+ (17); CCNH (21).
Repeated error correction, with higher-distance codes, has also been demonstrated. For single logical qubits, demonstrations include those on trapped-ion processors RBL+ (21); PBP+ (24) and superconducting processors KLR+ (22); Goo (23); GC (24); PNH+ (24) each with varying numbers of rounds, e.g., up to nearly $36$ rounds of error correction ($250$ syndrome extraction cycles) on the $\llbracket 49,1,7\rrbracket$ surface code, with a $0.143(3)\%$ logical error rate per cycle GC (24). Repeated error correction on multiple qubits has been implemented, including three rounds on two logical qubits PSR+ (24), and five rounds on eight logical qubits RAC+ (24), each while beating error rates of physical baselines.
II Physical machine description
Figure 3: Diagram of experimental architecture. 171Yb atoms, which each encode a physical qubit in their ground-state nuclear spin, can occupy tweezer traps within either a register array or an interaction zone (IZ). Up to 8 pairs of atoms interact at a time in the IZ, driven by two-qubit (2Q) gate lasers to implement controlled phase gates on the atoms in each pair. Arbitrary single-qubit (1Q) operations can be applied in parallel to the atoms within a row of the register; here one atom is missing. Each atom has a home register location, indicated by dashed lines, where it is originally loaded and where 1Q gates can be applied. Movement of atoms between the two zones enables arbitrary connectivity for 2Q gates. Up to eight atoms from the same row can move in parallel, but their paths cannot cross.
@@ II Physical machine description
All results presented in this work were obtained on a quantum processor based on reconfigurable arrays of neutral 171Yb atoms, depicted in Fig. 3, with the qubits encoded in the ground-nuclear spin states (1S0, $m_{F}=\pm 1/2$). Aspects of the system have previously been described in BBB+ (22); NC+ (23); NKC+ (24), as well as an accompanying publication MSS+ (24). Further system and benchmarking details will follow in a separate manuscript, in preparation.
By moving atoms between arrays of static optical tweezer traps using mobile tweezers, this platform enables arbitrary connectivity between qubits, for a large numbers of qubits. Single-qubit operations and qubit storage are performed in a register array, which is loaded in an iterative manner from a repeatedly filled reservoir NKC+ (24). Two-qubit gates are performed in a separate interaction zone (IZ). For this work, up to 256 atoms are available for computation, and up to eight atoms within the same row may be moved between the register and IZ at a time, subject to constraints associated with intermodulation in the acousto-optic deflectors used to generate the moving tweezers. Typical atom loss probabilities per move are below 0.1%. State-resolved, low-loss readout is performed by transferring atoms from the tweezer arrays into traps formed by cavity-enhanced optical lattices, which allows for state and atom detection with error and loss rates at the few parts-per-thousand level (see NKC+ (24) for a detailed characterization of state preparation and imaging). Importantly, the ability to perform state-selective, low-loss imaging allows us to distinguish between different types of error. An image performed prior to circuit execution allows the identification of sites successfully occupied by atoms, while a pair of images directly after circuit execution and after optical pumping within the qubit subspace allow determination of both the presence and state of the atoms.
In the encoded Bernstein-Vazirani experiment with $n=7$, for example, on average $0.52(7)\%$ of the $32$ atoms were not loaded, and $1054$ of $1200$ shots, or $85(2)\%$, had no initial loss. The fraction of qubits lost during the circuit averaged $2.3^{+0.2}_{-0.1}\%$. Initial loss varies, and more complicated circuits would have a higher circuit loss rate.
Single-qubit operations are performed in parallel on rows of atoms within the register, with the pulse area, phase, and presence of the gates controlled individually on each site using an architecture similar to that described in Ref. BBB+ (22). These gates are performed using two-photon Raman transitions driven by two independently controlled and site-resolved laser beams detuned roughly 10 GHz from the 3P1, $F=1/2$ manifold. We calibrate an $S_{X}$ gate on each site of the register, and perform virtual $Z$ rotations of arbitrary angle by changing the relative phase of the two drive lasers on subsequent gates. We characterize the performance of the single-qubit operations using a Clifford randomized benchmark sequence. At depths of up to 500 Clifford gates, we find an average fidelity of 99.85(2)per Clifford gate, averaged across three rows of the register spanning those used in this work. Because the nuclear spin qubit is well-isolated from differential light-shifts and has low sensitivity to magnetic fields, we do not automatically apply dynamical decoupling sequences for the circuits presented in this work.
Two-qubit controlled-phase (CZ) gates are performed in the IZ on up to eight pairs of atoms in parallel. These gates are mediated by excitation to a highly-excited Rydberg state ($65^{3}\mathrm{S}_{1}$). The sites of the IZ are spaced such that atoms occupying the same pair of sites are in the so-called “Rydberg blockade” limit, where the strength of interactions between the atoms exceeds other relevant energy scales, while atoms in different pairs interact weakly enough to be treated independently. All operations in the IZ are applied in parallel to all atoms within the IZ.
Atoms are excited to the Rydberg state using a sequential two-photon transition. State-selectivity is provided by the first stage of the two-photon transition, from 1S0, $m_{F}=-1/2$ to the metastable 3P0, $m_{F}=1/2$ state. Polarization selectivity, along with the narrow linewidth of the transition and slow Rabi rate (10 kHz) relative to the qubit frequency (376 kHz) ensure that only one qubit state is excited to the metastable state. From the metastable state, atoms are excited to the Rydberg state using a UV laser pulse (total duration 110 ns) that implements a gate similar to the time-optimal controlled-phase (CZ) gate described in JP (22). While the IZ traps are equally confining for the ground and metastable states, the register traps are near a tune-out wavelength for the metastable states, providing near-zero trapping potential. Thus, any gate errors that lead to population being left in the metastable state (such as some coherent population errors on either stage of the two-photon transition) are converted to loss and can be detected. This “erasure conversion” proves useful in identifying the location of errors WKPT (22).
We quantify the performance of our two-qubit gates using several variations on randomized benchmarking. In all presented cases, data is taken in parallel over the eight pairs of sites in the IZ, and averaged. To isolate the performance of the two-qubit gates themselves, in the absence of any errors that may be incurred through atomic movement, we perform interleaved randomized benchmarking (IRB) on static atoms within the IZ. In this protocol, we compare the return probability after a given number of random two-qubit Clifford operations (plus a return pulse calculated to return to a desired measurement state) to the same circuit with an additional CZ gate interleaved after each Clifford gate (and with an appropriately modified return pulse). By comparing the probability of returning to the measurement state versus the depth of the Clifford circuit with and without the added CZ gates, we isolate the error associated with the CZ gate alone, averaged over the two-qubit Hilbert space. Here, the atoms are kept in the IZ for all operations (unlike typical circuits, single qubit gates are applied in the IZ for this protocol). Fits to IRB data taken out to depth 34 indicate an infidelity per CZ gate of 0.56(15)and additional loss and population leakage outside of the qubit subspace of 0.27(6)%.
@@
Table 2: Randomized benchmark infidelities, post-selected on an atom (1Q) or pair of atoms (2Q) remaining in the qubit subspace, and probability of one or both atoms being lost from the qubit subspace. Benchmarking sequences are described in the main text.
@@
2Q Echoed RB with moves 0.4(1)0.5(1)%
@@
In actual circuits, atoms are moved between the register and IZ between single-qubit operations and two-qubit operations, to allow arbitrary connectivity. Because atom movement can cause heating, which may impact the performance of all following gates, IRB does not cleanly isolate the performance of the CZ gate when atoms are moved between arrays. For this, we rely on a variant of randomized benchmarking described in MSS+ (24) and similar to that in EBK+ (23), where concatenated circuit blocks each consist of a random single-qubit operation and two CZ gates separated by an echo pulse. A final gate sequence returns atoms to the measurement basis. We isolate the contribution from the two-qubit gates (along with any loss or infidelity due to associated movements, up to single-qubit phases which are canceled by the echo pulses) by subtracting the per-depth infidelity associated with the single-qubit operations. From this, we extract a per-CZ gate error of 0.4(1)and loss of 0.5(1)%. For comparison, the same procedure applied to atoms that remain static in the IZ for all operations yields a per-gate error rate of 0.35(5)and loss of 0.24(6)%. Table 2 summarizes the results of these benchmarks. The performance of our gates is sensitive to relative alignment of gate lasers and optical potentials, and other factors that may drift on the timescale of hours to days. The values presented here represent typical values over the time period in which the circuits presented in this work were run.
III Four-qubit codes
$\begin{array}[]{cccc}X&X&X&X\\ Z&Z&Z&Z\end{array}$
and logical operators
Observe that encoded ${|00\rangle}$ is the four-qubit cat state $\tfrac{1}{\sqrt{2}}({|0000\rangle}+{|1111\rangle})$. As shown in Figs. 4 and 5, an extra ancilla qubit is needed to prepare this state fault tolerantly Sho (96).
$S_{Z}XS_{Z}^{\dagger}$ $=Y$ $S_{Z}YS_{Z}^{\dagger}$ $=-X$ $S_{Z}ZS_{Z}^{\dagger}$ $=Z$
$S_{X}XS_{X}^{\dagger}$ $=X$ $S_{X}YS_{X}^{\dagger}$ $=Z$ $S_{X}ZS_{X}^{\dagger}$ $=-Y.$
IV Encoded $24$-qubit cat state
@@
An $n$-qubit cat state can be written as $\tfrac{1}{\sqrt{2}}({|0^{n}\rangle}+{|1^{n}\rangle})$. Here, we prepare an encoded cat state on $24$ logical qubits, encoded with the $\llbracket 4,2,2\rrbracket$ code into $48$ physical qubits. The circuit we implement is explained in Fig. 4. Starting with a logical circuit to prepare $\tfrac{1}{\sqrt{2}}({|0^{24}\rangle}+{|1^{24}\rangle})$, we fault-tolerantly encode it into the $\llbracket 4,2,2\rrbracket$ code. We then arrange the qubits, four code blocks in each row with the ancilla qubits at the end, and schedule the gates and qubit movements as in Fig. 5. The unencoded baseline experiment prepares a $24$-qubit cat state using a separately optimized circuit, in Fig. 6.

### v3

# Fault-tolerant quantum computation with a neutral atom processor

@@ Abstract
Quantum computing experiments are transitioning from running on physical qubits to using encoded, logical qubits. Fault-tolerant computation can identify and correct errors, and has the potential to enable the dramatically reduced logical error rates required for valuable algorithms. However, it requires flexible control of high-fidelity operations performed on large numbers of qubits. We demonstrate fault-tolerant quantum computation on a quantum processor with $256$ qubits, each an individual neutral Ytterbium atom. The operations are designed so that key error sources convert to atom loss, which can be detected by imaging. Full connectivity is enabled by atom movement. We demonstrate the entanglement of $24$ logical qubits encoded into $48$ atoms, at once catching errors and correcting for, on average 1.8, lost atoms. We also implement the Bernstein-Vazirani algorithm with up to $28$ logical qubits encoded into $112$ atoms, showing better-than-physical error rates. In both cases, “erasure conversion,” changing errors into a form that can be detected independently from qubit state, improves circuit performance. These results begin to clear a path for achieving scientific quantum advantage with a programmable neutral atom quantum processor.
Quantum computers are potentially powerful, but are challenging to control precisely. State-of-the-art two-qubit gate error rates in some flagship quantum processors range from $0.1\%$ for a trapped-ion device MBA+ (23) to an average of $0.4\%$ for a superconducting qubit device GC (24). Moderate-size quantum algorithms require much lower error rates, below $10^{-6}$ BMT+ (22). This precision can likely only be achieved using quantum error-correcting codes to store and fault-tolerantly operate on “logical” qubits AGP (06). Here, we advance the shift from physical to logical qubits, by demonstrating fault-tolerant, encoded computations on logical qubits. The encoded circuits outperform unencoded circuits on physical qubits.
Our results are achieved using a neutral atom quantum processor with $256$ qubits, Ytterbium atoms (Fig. 1), designed to enable loss conversion and detection for improved logical performance. Neutral atom devices have been mainly analog, but recent progress has enabled digital, gate-based computation EBK+ (23); SST+ (23); MLP+ (23); RCC+ (24); FTS+ (24); MSS+ (24). The processor here has low gate error, and atom movement enables all-to-all connectivity for flexible computation and logical operation design. The physical two-qubit gates are designed so that key error sources convert to atom loss, which imaging has been designed to detect. With unencoded circuits, data with lost atoms can be discarded, and using this selection we prepare an entangled $40$-qubit “cat” state (Fig. 2). Loss conversion is even more advantageous operating on encoded data, though; a distance-two code can detect errors and correct for lost atoms. We demonstrate the creation and entanglement of logical qubits, and logical computation. We prepare a cat state on $24$ logical qubits, encoded in $48$ atoms (Fig. 3). We then demonstrate a quantum algorithm on logical qubits. We fault-tolerantly run the Bernstein-Vazirani algorithm BV (97); CEMM (98), well known for showing the higher efficiency of quantum computers versus classical, on as many as $28$ logical qubits, encoded in $112$ atoms (Fig. 4), with better-than-physical results. This illustrates the scale available in neutral atom processors, and their promise as reliable, fault-tolerant quantum processors.
Figure 1: Experimental architecture. a, b, 171Yb atoms, each giving a physical qubit in the ground-state nuclear spin, can occupy tweezer traps within either a register array or an interaction zone (IZ). Up to 8 pairs of atoms interact at a time in the IZ, driven by two-qubit (2Q) gate lasers to implement controlled-phase (CZ) gates. These lasers perform coherent sequential excitation to a metastable clock state (3P0) and the Rydberg state (653S1). Arbitrary single-qubit (1Q) operations can be applied in parallel to the atoms within a register row using Raman transitions detuned from the 3P1 manifold. Moving atoms between the zones enables arbitrary 2Q gate connectivity. c, Circuit-running sequence. Dissipative operations (left column) are performed with atoms held in a cavity-enhanced optical lattice. Three images (orange boxes) allow us to determine initial occupancy of register sites, as well as the final state and presence of atoms. Coherent operations (right column) are performed with atoms held in tweezers.
Figure 2: Fidelity of prepared cat states versus qubit number $n$. Red points give fidelities for trials with no lost atoms. Cat states with up to $40$ qubits exceed the entanglement threshold. Gray points include trials in which all atoms are present initially but atoms lost during the circuit are replaced by the maximally mixed state, and black points include all trials.
Figure 3: Encoded $24$-qubit cat state results. Shown is the error for measurements in the $X$ basis, the error for measurements in the $Z$ basis, and their sum, for three cases. The baseline is an unencoded $24$-qubit cat state, based on trials with no lost atoms. For the encoded cat state data, in one case we reject trials with a detected error, but try to correct for qubits lost during the circuit. In the other case, we reject trials with any lost qubit or detected error.
@@
Figure 4: Results for the Bernstein-Vazirani algorithm with an $n$-bit secret string. The string is always $1^{n}$, as the corresponding oracle is the most difficult to implement. The success probability is the probability of measuring $1^{n}$, conditioned on acceptance. For the unencoded algorithm (black), we accept when the $n$ measured atoms are present. For the encoded algorithm (red), we accept when all of the $n$ measured code blocks are decodable; in particular, it can detect faults and correct for lost atoms. The lighter points are based only on those trials with no initial atom loss.
Creating and verifying large cat or GHZ states GHZ (89) is a common benchmark for hardware progress Kre (24). The current record is a $60$-qubit cat state, prepared with $59\%$ fidelity using superconducting qubits BXS+ (24). A $20$-qubit cat state with $54\%$ fidelity, after adjusting for measurement error, has been prepared on a neutral atom system OLK+ (19). The $62^{+5}_{-6}\%$ fidelity of our $40$-qubit cat state is not adjusted for measurement errors.
@@
Table 1 summarizes results with encoded cat states. A cat state on $12$ logical qubits has been prepared with $99.8\%$ fidelity RAC+ (24), on a trapped-ion processor MBA+ (23). On a neutral atom system, Bluvstein et al. BEG+ (24) have prepared a four-qubit encoded cat state with fidelity $72(2)\%$ using error correction, ranging to $99.85^{+0.1}_{-1.0}\%$ using full error detection and with a lower acceptance rate. Our $24$-qubit encoded cat state is the largest published to date.
@@
Reference qubits Code Total error $p_{X}+p_{Z}$
HDHL (24) 4 $\llbracket 25,4,3\rrbracket$ $0.5^{+0.4}_{-0.2}\%$
@@
28(2)error correction
$0.15^{+1.0}_{-0.1}\%$ error detection
RAC+ (24) 12 $\llbracket 16,4,4\rrbracket$ $0.18^{+0.4}_{-0.12}\%$
@@
$10.2^{+6.6}_{-4.7}\%$ error and
loss detection
@@
The Bernstein-Vazirani algorithm for finding a secret string is another benchmark for quantum processors LGM+ (24). It has been implemented with unencoded qubits on superconducting processors RHK+ (20); DH (21); PL (23), ion trap processors FHM+ (16); DLF+ (16); WBD+ (19), and both processor types LMRM (17); MLM+ (19); BWM (21); LJV+ (23), for various lengths of the secret string, up to length $26$ for superconducting qubits PL (23) and up to length $20$ for ion traps LJV+ (23). It is difficult to compare these results fairly. Some test identification of the secret $s=1^{n}$, since it is hardest to implement; the corresponding oracle requires the most two-qubit gates. Others test against random strings $s$, and this is much easier, comparable to finding $s=1^{n/2}$. The success probabilities also vary considerably, e.g., from $0.2\%$ for $s=1^{26}$ PL (23), to $70\%$ for random $s$ with $n=20$ LJV+ (23). Our fault-tolerant, encoded implementation of the Bernstein-Vazirani algorithm is the first to date, and has the largest length, $n=27$. On $s=1^{27}$, the $91.4(3)\%$ success probability for the encoded circuit beats the unencoded algorithm.
Neutral atom quantum processor
@@
The quantum processor used in this work is based on reconfigurable arrays of neutral 171Yb atoms, depicted in Fig. 1a,b, with the qubits encoded in the ground-nuclear spin states (1S0, $m_{F}=\pm 1/2$). Aspects of the system have been described in previous publications BBB+ (22); NC+ (23); NKC+ (24); MSS+ (24), and further details can be found in the methods section. By moving atoms between arrays of static optical tweezer traps using mobile tweezers, this platform enables arbitrary connectivity between large numbers of qubits.
Figure 1c shows a typical sequence for circuit execution. State preparation consists of iterative loading of the relevant sites of the register array, as described in NKC+ (24), followed by an image to determine successful filling of sites, cooling, and initialization of the ${|0\rangle}$ state through optical pumping. State-resolved, low-loss readout as well as cooling and state initialization are performed by transferring atoms from the tweezer arrays into traps formed by cavity-enhanced optical lattices, which allows for state and atom detection with error and loss rates at the part-per-thousand level (see methods). All other operations are peformed with the atoms held in the tweezer arrays. For this work, up to 256 atoms are available for computation.
During circuit execution, single-qubit operations and qubit storage are performed in a register array. Two-qubit (2Q) gates are performed on up to eight qubit pairs in parallel within a separate interaction zone (IZ), with connectivity determined by the placement of atoms within the IZ. 2Q gates are performed using state-selective sequential two-photon excitation to a high-lying Rydberg state in a blockaded regime, resulting in controlled-phase (CZ) gates between pairs of qubits. Importantly, key sources of 2Q gate errors are converted to loss, which can be detected independently from the qubit state. Population errors in the sequential excitation used in the gates lead either directly to loss, or to population of metastable states that are not trapped in register traps. Unlike most systems based on individually controlled neutral atoms, our ability to perform nondestructive state-selective imaging NC+ (23) allows us to perform a pair of images that unambiguously distinguish between three possibilities for the atomic state: ${|0\rangle}$, ${|1\rangle}$, and lost.
Table 2 presents a summary of important error rates, with benchmark protocols described in the methods.
@@
Table 2: Randomized benchmark (RB) infidelities, conditioned on an atom (1Q) or pair of atoms (2Q) staying in the qubit subspace, and probability of losing one or both atoms.
@@
2Q Echoed RB with moves 0.39(10)0.75(10)%
@@
Cat state generation
We prepare non-encoded cat states using modified versions of the circuits from CFG+ (19). We generate the “antiferromagnetic” cat state $\frac{1}{\sqrt{2}}({|0101\ldots\rangle}+{|1010\ldots\rangle})$, instead of the “ferromagnetic” state $\frac{1}{\sqrt{2}}({|0^{n}\rangle}+{|1^{n}\rangle})$, since it is less sensitive to correlated phase errors. Figure 2 plots the fidelity to the ideal cat state, computed as $1-\tfrac{1}{2}(Z$ basis error rate)()-() - (average error rate over $n$ bases orthogonal to $z)$ GLGP (07). A fidelity above 50% certifies $n$-partite entanglement GS (10). Restricting the analysis to only those trials where all atoms survived, the cat states exceed the entanglement threshold up to $n=40$. To our knowledge, this is the largest cat state demonstrated with neutral atoms.
We can also estimate the fidelity including all trials, or all trials with no initial loss, in the analysis generously treating lost qubits as replaced by the maximally mixed state. This gives a higher acceptance rate, but lower fidelity, showing the advantage of loss detection.
Encoded $24$-qubit cat state
@@
We prepare a cat state on $24$ logical qubits, the largest published to date. The logical qubits are encoded with the $\llbracket 4,2,2\rrbracket$ code into $48$ physical qubits. Starting with a logical state-preparation circuit, we encode the circuit, then schedule the gates and movements (Fig. 5).
Figure 5: a, Logical circuit to prepare a $24$-qubit cat state. The qubits are arranged in $12$ pairs that will each be encoded into a $\llbracket 4,2,2\rrbracket$ code block. b, The encoded circuit involves $59$ atoms and $101$ physical CZ gates. All but the first code block use extra ancilla qubits, shown in gray, to prepare encoded ${|00\rangle}$ fault tolerantly; each ancilla should be measured to be 00. In the diagram, a blue wire is a qubit in the Hadamard basis. Not shown are the final transversal $X$ or $Z$ measurements. c, Schedule for the atom movements and gates.
Our experiments, like HDHL (24); RAC+ (24), measure $X$ basis and $Z$ basis error rates, $p_{X}$ and $p_{Z}$. In the unencoded case, the fidelity to the ideal cat state is at least $1-p_{X}-p_{Z}$. We find that for the encoded cat state, using error detection and loss correction, $p_{X}+p_{Z}=26.6^{+3.3}_{-3.1}\%$, while for the unencoded baseline $p_{X}+p_{Z}=42.0^{+11.3}_{-10.1}\%$. The results are summarized in Fig. 3. The baseline error bars are large because we report the best single run (different from in Fig. 2), which disadvantages the encoded cases. The encoding is beneficial, but affects $p_{X}$ and $p_{Z}$ differently.
How useful is loss conversion, loss measurement, and loss correction for this encoded state? The average loading loss rate was $1.04(2)\%$, and $57.2(6)\%$ of trials had no initially lost atoms. Among trials without initial losses, those that decoded correctly had an average of 2.0 and 1.6 circuit losses, for $X$ and $Z$ basis measurements, respectively. The code successfully corrects for these lost qubits. Still, circuit loss is a major cause of logical errors. Those shots that decoded to a logical error had respective averages of 2.4 and 2.7 losses. In the $Z$ basis, there are only two valid decodings, $0^{24}$ and $1^{24}$, and we can usually identify the most likely cause of a logical error. For example, a decoding of $1111(0111)^{5}$ indicates that most probably the third block had a logical $XI$ error that propagated to the subsequent code blocks. The frequencies of the most likely causes of the logical error are: $26\%$ two circuit losses, $54\%$ one loss and one $X$ error, $9\%$ two $X$ errors, $11\%$ three or more losses or errors.
Figure 6 shows how conditioning on at most $k$ lost qubits, for $k=0,1,\ldots$, allows an interpolation between the different encoded results in Fig. 3; in particular, conditioning on no loss reduces the total error from $29.5(3.0)\%$ to $10.2^{+6.6}_{-4.7}\%$. Finally, Fig. 7 shows all the encoded cat state runs, plotted conditioned on no initial atom loss. The spread and correlations in the figure suggest some machine parameter drift.
Figure 6: Encoded $24$-qubit cat state data, selecting only trials with specified loss criteria. Trying to correct more loss generally leads to a higher acceptance rate, but also a higher error rate. In each trial, an atom may be missing before the circuit begins or it may be lost during the circuit. The red points show the total $X$ basis plus $Z$ basis error rate, plotted against the average of the $X$ and $Z$ acceptance rates, based on either all 29 7602976029\,76029 760 trials, only those 17 0081700817\,00817 008 trials with no initial loss, or only those $858$ trials with no loss at all. The blue points select only trials in which the total number of lost qubits is at most a threshold.
Figure 7: Encoded $24$-qubit cat state runs. Each point gives the acceptance probability and logical error rate for a run of, on average, $183$ trials with no initial loss, blue for the $X$ basis and red for $Z$. For clarity, only some error bars are shown. The large squares are the full averages, as reported in Fig. 3. The negative correlation between acceptance probability and logical error rate evidences some parameter drift between runs, despite automatic recalibrations.
Encoded Bernstein-Vazirani algorithm
Computing with logical qubits is the next step after encoded state preparation. For a secret string $s\in\{0,1\}^{n}$, consider a black-box oracle that on query $x\in\{0,1\}^{n}$ returns $s\cdot x=s_{1}x_{1}\oplus\cdots\oplus s_{n}x_{n}$. Classically, $n$ oracle queries are required to learn $s$. The Bernstein-Vazirani (BV) algorithm BV (97); CEMM (98) makes a single call to the corresponding quantum oracle $U_{s}$, $U_{s}{|b,x\rangle}={|b\oplus(s\cdot x),x\rangle}$, to learn $s$ via $H^{n+1}U_{s}H^{n+1}{|1,0^{n}\rangle}={|1,s\rangle}$. This algorithm was designed to prove an oracle separation between the classical randomized and quantum polynomial-time complexity classes $\mathsf{BPP}$ and $\mathsf{BQP}$.